# DualNILM: Energy Injection Identification enabled Disaggregation with Deep Multi-Task Learning

### v2

@@ V-B Evaluation Metrics
For the energy injection estimation task, we use the Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) as the metrics:
@@
V-C Benchmark Methods
To comprehensively evaluate DualNILM, we compare it with representative and state-of-the-art methods in both appliance state recognition and energy injection disaggregation tasks. Table II summarizes our benchmark methods.
| Category | Method | Key Characteristics | Primary Strengths | Tasks* | ||
| Traditional | FHMM [7] | Parallel hidden Markov chains | Interpretable state modeling | Both | ||
| XGBoost [41] | Tree-based ensemble | Non-linear pattern recognition | State Recognition | |||
Seq2Point [11] | CNN-based architecture | Local temporal features | State Recognition | ||
| CNN-LSTM [42] | Hybrid CNN-LSTM | Local-global temporal modeling | State Recognition | |||
| Transformer [27] | Self-attention mechanism | Long-range dependencies | State Recognition | |||
| Seq2Seq [13] | Encoder-decoder LSTM | Sequence mapping | Disaggregation | |||
| DAE [43] | Denoising autoencoder | Signal reconstruction | Disaggregation | |||
| UNetNILM [44] | Skip-connected encoder-decoder | Multi-task learning | Both | |||
| *Tasks include State Recognition (appliance state recognition), Disaggregation (energy injection disaggregation), and Both (both tasks). | ||||||
For fair comparison and evaluation, we select both traditional NILM methods, such as FHMM, and modern deep learning approaches. Although some methods specialize in appliance state recognition or energy injection disaggregation, methods such as FHMM and UNetNILM handle both tasks simultaneously. This diverse selection enables comprehensive evaluation of DualNILM’s performance across different technical approaches and computational paradigms, with particular attention to the unique challenges posed by BTM energy injection.
Detailed configurations and implementation specifics are provided in Appendix D-A. All benchmark models were implemented with carefully tuned hyperparameters to ensure fair comparison. This comprehensive selection enables evaluation of different technical approaches, with particular attention to methods capable of handling BTM energy injection.
VI Experimental Results
@@
Accuracy (%) | Recall (%) | Precision (%) | F1 (%) | |||
| Laboratory Dataset (Real Injection) | FHMM | 60.95 | 28.32 | 16.84 | 17.25 | |||
| XGBoost | 85.82 | 48.20 | 52.10 | 47.75 | ||||
| Seq2Point | 67.63 | 53.06 | 27.86 | 34.27 | ||||
| CNN-LSTM | 74.03 | 52.97 | 30.29 | 36.67 | ||||
| Transformer | 81.79 | 52.91 | 38.16 | 42.76 | ||||
| UNetNILM | 89.20 | 54.01 | 60.81 | 53.89 | ||||
| DualNILM | 98.54 | 85.43 | 84.16 | 84.54 | ||||
| REDD with Synthetic Injection | House 1 | FHMM | 79.52 | 1.46 | 13.47 | 2.07 | ||
| XGBoost | 97.90 | 74.50 | 88.74 | 79.18 | ||||
| Seq2Point | 97.65 | 42.20 | 56.42 | 46.07 | ||||
| CNN-LSTM | 98.67 | 46.32 | 50.75 | 47.96 | ||||
| Transformer | 95.99 | 56.08 | 54.10 | 53.42 | ||||
| UNetNILM | 99.05 | 76.11 | 85.22 | 75.27 | ||||
| DualNILM | 99.84 | 96.26 | 97.11 | 96.64 | ||||
| House 2 | FHMM | 75.08 | 0.58 | 0.34 | 0.43 | |||
| XGBoost | 98.43 | 92.47 | 95.58 | 93.53 | ||||
| Seq2Point | 91.43 | 52.90 | 58.44 | 50.85 | ||||
| CNN-LSTM | 98.86 | 49.19 | 48.74 | 48.96 | ||||
| Transformer | 97.92 | 47.78 | 47.98 | 47.87 | ||||
| UNetNILM | 99.57 | 64.16 | 77.22 | 67.28 | ||||
| DualNILM | 99.70 | 92.48 | 94.00 | 93.11 | ||||
| House 3 | FHMM | 83.84 | 4.53 | 0.90 | 1.49 | |||
| XGBoost | 94.89 | 46.22 | 56.38 | 50.51 | ||||
| Seq2Point | 98.10 | 42.78 | 50.90 | 44.47 | ||||
| CNN-LSTM | 95.92 | 28.85 | 38.17 | 32.62 | ||||
| Transformer | 97.81 | 40.21 | 38.89 | 39.43 | ||||
| UNetNILM | 99.07 | 59.91 | 54.73 | 57.10 | ||||
| DualNILM | 99.69 | 77.42 | 84.46 | 80.04 | ||||
| UKDALE with Synthetic Injection | House 1 | FHMM | 89.25 | 1.58 | 0.76 | 1.01 | ||
| XGBoost | 93.64 | 41.74 | 68.04 | 45.97 | ||||
| Seq2Point | 88.70 | 41.47 | 43.29 | 36.79 | ||||
| CNN-LSTM | 98.33 | 31.76 | 31.29 | 31.52 | ||||
| Transformer | 96.90 | 33.65 | 27.76 | 30.26 | ||||
| UNetNILM | 98.76 | 52.03 | 72.01 | 55.20 | ||||
| DualNILM | 99.72 | 95.23 | 94.30 | 94.71 | ||||
| House 2 | FHMM | 82.95 | 0.62 | 0.52 | 0.55 | |||
| XGBoost | 96.13 | 66.61 | 79.19 | 70.67 | ||||
| Seq2Point | 98.83 | 42.68 | 46.49 | 41.92 | ||||
| CNN-LSTM | 97.81 | 41.65 | 45.48 | 43.33 | ||||
| Transformer | 99.13 | 59.88 | 77.44 | 61.73 | ||||
| UNetNILM | 99.17 | 64.25 | 68.74 | 65.43 | ||||
| DualNILM | 99.73 | 90.63 | 94.46 | 92.31 | ||||
@@
Tables III and IV report comprehensive performance metrics for appliance-state recognition and energy-injection disaggregation. We evaluated them in two complementary settings: i) controlled laboratory experiments using a real 500 W micro-inverter as the BTM injection; and ii) household datasets (REDD and UK-DALE) augmented with synthetic PV profiles to emulate typical residential consumption–generation patterns. This multi-faceted design enables robust validation across diverse, realistic household usage regimes.
VI-A Performance on Laboratory Dataset
@@
The laboratory experiments provide critical insights into real-world performance with actual hardware effects. Table III (first block) reveals fundamental challenges that bidirectional power flows pose to conventional NILM methods. FHMM, despite its theoretical ability to model multiple hidden states, achieves only 17.25% F1-score. This poor performance stems from its factorial hidden Markov model structure, which assumes discrete state transitions and cannot accommodate the continuous power variations introduced by inverter operation. The model essentially interprets injection-induced fluctuations as rapid appliance state changes, leading to excessive false positives.
@@
For continuous injection estimation (Table IV), the laboratory results reveal interesting trade-offs between different regression approaches. Seq2Seq achieves competitive MAE (0.2622) through its sequence-to-sequence architecture optimized for time series prediction. However, its higher RMSE (0.3512) indicates susceptibility to outliers during rapid injection transitions. DAE shows the poorest performance (RMSE: 0.8830), as its denoising objective is misaligned with the injection estimation task. UNetNILM demonstrates reasonable performance (RMSE: 0.2305) by leveraging its multiscale feature extraction.
@@
| Laboratory Dataset (Real Injection) | FHMM | 0.5015 | 0.4492 | |||||
| Seq2Seq | 0.3512 | 0.2622 | ||||||
| DAE | 0.8830 | 0.7851 | ||||||
| UNetNILM | 0.2305 | 0.1883 | ||||||
| DualNILM | 0.1429 | 0.0804 | ||||||
| REDD with Synthetic Injection | House 1 | FHMM | 0.8093 | 0.2200 | ||||
| Seq2Seq | 0.6834 | 0.2370 | ||||||
| DAE | 0.8115 | 0.2171 | ||||||
| UNetNILM | 0.7589 | 0.1983 | ||||||
| DualNILM | 0.6661 | 0.1777 | ||||||
| House 2 | FHMM | 0.8403 | 0.3367 | |||||
| Seq2Seq | 0.7217 | 0.3287 | ||||||
| DAE | 0.8448 | 0.3356 | ||||||
| UNetNILM | 2.2077 | 0.6601 | ||||||
| DualNILM | 0.7065 | 0.2768 | ||||||
| House 3 | FHMM | 0.5575 | 0.1766 | |||||
| Seq2Seq | 0.4974 | 0.1940 | ||||||
| DAE | 0.5606 | 0.1743 | ||||||
| UNetNILM | 2.2517 | 0.4708 | ||||||
| DualNILM | 0.4261 | 0.1422 | ||||||
| UKDALE with Synthetic Injection | House 1 | FHMM | 1.1905 | 0.3331 | ||||
| Seq2Seq | 0.9055 | 0.2758 | ||||||
| DAE | 0.9197 | 0.2948 | ||||||
| UNetNILM | 7.5014 | 1.3871 | ||||||
| DualNILM | 0.9008 | 0.2731 | ||||||
| House 2 | FHMM | 1.1477 | 0.3521 | |||||
| Seq2Seq | 0.9252 | 0.2833 | ||||||
| DAE | 0.9594 | 0.3368 | ||||||
| UNetNILM | 1.1285 | 0.3043 | ||||||
| DualNILM | 0.9344 | 0.2883 | ||||||
VI-B Performance on Synthetic Public Datasets
@@
Table IV reveals that injection estimation becomes more challenging on public datasets compared to laboratory settings. This is expected as the synthetic injection exhibits more complex temporal dynamics including cloud transients and seasonal variations. Seq2Seq occasionally achieves competitive RMSE on individual houses but shows less consistent MAE performance. Notably, UNetNILM exhibits severe instability on UK-DALE House 1 (RMSE: 7.5014), suggesting training convergence issues when the injection magnitude varies widely. This instability, contrasted with its reasonable performance on REDD datasets, highlights the sensitivity of standard multi-task architectures to dataset-specific characteristics without proper architectural design for handling bidirectional power flows.
DualNILM delivers balanced performance across all metrics and houses. While not always achieving the absolute lowest RMSE on every house, it consistently ranks among the top performers for both RMSE and MAE, demonstrating robust generalization. The multi-task learning enables better handling of the complex interaction between varying consumption patterns and injection profiles.
VI-C Key Insights and Analysis
@@
Traditional methods, such as FHMM, fail due to violated mathematical assumptions, and their factorial structure assumes discrete states incompatible with continuous injection variations. Notably, XGBoost serves as a strong baseline that achieves competitive performance (74.41% average F1 on REDD from Table XIII) through implicit feature engineering via recursive tree splits on the dual-channel input. However, it is fundamentally limited by its inability to provide explicit injection estimates and its reliance on pattern memorization, which makes it susceptible to distributional shifts.
@@ VII Conclusion
Our findings underscore DualNILM’s potential for robust NILM in modern energy systems that integrate renewable energy sources, and our provided open research infrastructure will help the NILM community continue to advance research in this critical direction.
@@ Appendix A Theoretical Analysis of the Challenge: When Negative Flows Disrupt Classical NILM
A-A Matrix Formulation of NILM
@@
A-B Why BTM Injection Complicates NILM
In the language of signal processing and blind source separation (BSS), classical NILM can be viewed as a single-channel mixture of purely nonnegative sources. The geometry of such mixtures is conic: if each device’s load lies in the nonnegative orthant, standard factorization approaches (e.g., nonnegative matrix factorization) can leverage convexity to help isolate distinct appliance patterns [45].
@@ A-B Why BTM Injection Complicates NILM
A-C Relaxation to Negative Flows and Identifiability
@@
Degraded Identifiability and Solution Stability. Under classical NILM assumptions (fully nonnegative), more stable decompositions can sometimes be achieved when the underlying sources have distinct support or exhibit certain ”sparse plus low-rank” structures. These identifiability arguments are often reminiscent of theorems in blind source separation (BSS) and NMF, where the nonnegative cone enforces a geometric structure that ”pins down” solutions. Once negative values are allowed (even for a single source), the problem becomes closer to an unconstrained BSS, for which solution identifiability is well known to deteriorate significantly. In more formal terms, NMF-based identifiability relies on the fact that data points lie in a convex cone generated by the basis columns [45]; introducing negative components breaks those convex cone arguments.
A-D Signal Space Interpretation and Rank Constraints

### v3

@@ 5.2. Evaluation Metrics
For the BTM energy injection estimation task, we use the Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) as the metrics:
@@
5.3. Benchmark Methods
To comprehensively evaluate DualNILM, we compare it with representative and state-of-the-art methods in both appliance state recognition and energy injection disaggregation tasks. Table 2 summarizes our benchmark methods.
| Category | Method | Key Characteristics | Primary Strengths | Tasks* |
| Traditional | FHMM (Kim et al., 2011) | Parallel hidden Markov chains | Interpretable state modeling | Both |
| XGBoost (Chen and Guestrin, 2016) | Tree-based ensemble | Non-linear pattern recognition | State Recognition | |
| SunDance (Chen and Irwin, 2017) | Physical solar modeling | Black-box estimation | Disaggregation | |
| Deep Learning | Seq2Point (Zhang et al., 2018) | CNN-based architecture | Local temporal features | State Recognition |
| CNN-LSTM (Kaselimi et al., 2020) | Hybrid CNN-LSTM | Local-global temporal modeling | State Recognition | |
| Transformer (Vaswani et al., 2017) | Self-attention mechanism | Long-range dependencies | State Recognition | |
| Seq2Seq (Du et al., 2016) | Encoder-decoder LSTM | Sequence mapping | Disaggregation | |
| DAE (Jia et al., 2019) | Denoising autoencoder | Signal reconstruction | Disaggregation | |
| BERT+ (Li et al., 2024) | Bert Blocks | Ultilize extra user embeddings | Disaggregation | |
| UNetNILM (Faustine et al., 2020) | Skip-connected encoder-decoder | Multi-task learning | Both | |
| *Tasks include State Recognition (appliance state recognition), Disaggregation (energy injection disaggregation), and Both (both tasks). | ||||
For fair comparison and evaluation, we select both traditional NILM methods, such as FHMM, and modern deep learning approaches. Although some methods specialize in appliance state recognition or energy injection disaggregation, methods such as FHMM and UNetNILM handle both tasks simultaneously. We additionally benchmark against recent BTM-specific approaches including SunDance (Chen and Irwin, 2017) for solar disaggregation and BERT+ (Li et al., 2024) for injection estimation. This diverse selection enables comprehensive evaluation of DualNILM’s performance across different technical approaches and computational paradigms, with particular attention to the unique challenges posed by BTM energy injection.
Detailed configurations and implementation specifics are provided in Appendix D.1. All benchmark models were implemented with carefully tuned hyperparameters to ensure fair comparison. This comprehensive selection enables evaluation of different technical approaches, with particular attention to methods capable of handling BTM energy injection.
6. Experimental Results
@@
| Dataset | Method | Accuracy (%) | Recall (%) | Precision (%) | F1 (%) | |
| Laboratory Dataset (Real Injection) | FHMM | 60.95 | 28.32 | 16.84 | 17.25 | |
| XGBoost | 85.82 | 48.20 | 52.10 | 47.75 | ||
| Seq2Point | 67.63 | 53.06 | 27.86 | 34.27 | ||
| CNN-LSTM | 74.03 | 52.97 | 30.29 | 36.67 | ||
| Transformer | 81.79 | 52.91 | 38.16 | 42.76 | ||
| UNetNILM | 89.20 | 54.01 | 60.81 | 53.89 | ||
| DualNILM | 98.54 | 85.43 | 84.16 | 84.54 | ||
| REDD with Synthetic Injection | House 1 | FHMM | 79.52 | 1.46 | 13.47 | 2.07 |
| XGBoost | 97.90 | 74.50 | 88.74 | 79.18 | ||
| Seq2Point | 97.65 | 42.20 | 56.42 | 46.07 | ||
| CNN-LSTM | 98.67 | 46.32 | 50.75 | 47.96 | ||
| Transformer | 95.99 | 56.08 | 54.10 | 53.42 | ||
| UNetNILM | 99.05 | 76.11 | 85.22 | 75.27 | ||
| DualNILM | 99.84 | 96.26 | 97.11 | 96.64 | ||
| House 2 | FHMM | 75.08 | 0.58 | 0.34 | 0.43 | |
| XGBoost | 98.43 | 92.47 | 95.58 | 93.53 | ||
| Seq2Point | 91.43 | 52.90 | 58.44 | 50.85 | ||
| CNN-LSTM | 98.86 | 49.19 | 48.74 | 48.96 | ||
| Transformer | 97.92 | 47.78 | 47.98 | 47.87 | ||
| UNetNILM | 99.57 | 64.16 | 77.22 | 67.28 | ||
| DualNILM | 99.70 | 92.48 | 94.00 | 93.11 | ||
| House 3 | FHMM | 83.84 | 4.53 | 0.90 | 1.49 | |
| XGBoost | 94.89 | 46.22 | 56.38 | 50.51 | ||
| Seq2Point | 98.10 | 42.78 | 50.90 | 44.47 | ||
| CNN-LSTM | 95.92 | 28.85 | 38.17 | 32.62 | ||
| Transformer | 97.81 | 40.21 | 38.89 | 39.43 | ||
| UNetNILM | 99.07 | 59.91 | 54.73 | 57.10 | ||
| DualNILM | 99.69 | 77.42 | 84.46 | 80.04 | ||
| UKDALE with Synthetic Injection | House 1 | FHMM | 89.25 | 1.58 | 0.76 | 1.01 |
| XGBoost | 93.64 | 41.74 | 68.04 | 45.97 | ||
| Seq2Point | 88.70 | 41.47 | 43.29 | 36.79 | ||
| CNN-LSTM | 98.33 | 31.76 | 31.29 | 31.52 | ||
| Transformer | 96.90 | 33.65 | 27.76 | 30.26 | ||
| UNetNILM | 98.76 | 52.03 | 72.01 | 55.20 | ||
| DualNILM | 99.72 | 95.23 | 94.30 | 94.71 | ||
| House 2 | FHMM | 82.95 | 0.62 | 0.52 | 0.55 | |
| XGBoost | 96.13 | 66.61 | 79.19 | 70.67 | ||
| Seq2Point | 98.83 | 42.68 | 46.49 | 41.92 | ||
| CNN-LSTM | 97.81 | 41.65 | 45.48 | 43.33 | ||
| Transformer | 99.13 | 59.88 | 77.44 | 61.73 | ||
| UNetNILM | 99.17 | 64.25 | 68.74 | 65.43 | ||
| DualNILM | 99.73 | 90.63 | 94.46 | 92.31 | ||
@@
Tables 3 and 7 report comprehensive performance metrics for appliance-state recognition and energy-injection disaggregation. We evaluated them in two complementary settings: i) controlled laboratory experiments using a real 500 W micro-inverter as the BTM injection; and ii) household datasets (REDD and UK-DALE) augmented with synthetic PV profiles to emulate typical residential consumption–generation patterns. This multi-faceted design enables robust validation across diverse, realistic household usage regimes. Figure 5 shows the Avg. F1-score performance comparison across all datasets.
6.1. Performance on Laboratory Dataset
@@
The laboratory experiments provide critical insights into real-world performance with actual hardware effects. Table 3 (first block) reveals fundamental challenges that bidirectional power flows pose to conventional NILM methods. FHMM, despite its theoretical ability to model multiple hidden states, achieves only 17.25% F1-score. This poor performance stems from its factorial hidden Markov model structure, which assumes discrete state transitions and cannot accommodate the continuous power variations introduced by inverter operation. The model essentially interprets injection-induced fluctuations as rapid appliance state changes, leading to excessive false positives.
@@
For continuous injection estimation (Figure 6 and Appendix Table 7), the laboratory results reveal interesting trade-offs between different regression approaches. Seq2Seq achieves competitive MAE (0.2622) through its sequence-to-sequence architecture optimized for time series prediction. However, its higher RMSE (0.3512) indicates susceptibility to outliers during rapid injection transitions. DAE shows the poorest performance (RMSE: 0.8830), as its denoising objective is misaligned with the injection estimation task. UNetNILM demonstrates reasonable performance (RMSE: 0.2305) by leveraging its multiscale feature extraction.
@@
6.2. Performance on Synthetic Public Datasets
@@
Figure 6 and Table 7 in Appendix reveal that injection estimation becomes more challenging on public datasets compared to laboratory settings. This is expected as the synthetic injection exhibits more complex temporal dynamics including cloud transients and seasonal variations. Seq2Seq occasionally achieves competitive RMSE on individual houses but shows less consistent MAE performance. SunDance (Chen and Irwin, 2017), designed specifically for solar disaggregation using physical generation models, exhibits limited performance compared with other NILM disaggregation baselines, while BERT+ demonstrates competitive performance (REDD: 0.6590 average RMSE; UK-DALE: 0.9919 average RMSE). Notably, UNetNILM exhibits severe instability on UK-DALE House 1 (RMSE: 7.5014 in full Table 7), suggesting training convergence issues when the injection magnitude varies widely. This instability, contrasted with its reasonable performance on REDD datasets, highlights the sensitivity of standard multi-task architectures to dataset-specific characteristics without proper architectural design for handling bidirectional power flows. DualNILM delivers balanced performance across all metrics and houses. While not always achieving the absolute lowest RMSE on every house, it consistently ranks among the top performers for both RMSE and MAE, demonstrating robust generalization. The multi-task learning enables better handling of the complex interaction between varying consumption patterns and injection profiles.
6.3. Ablation Study on DualNILM
To validate our architectural design choices, we conducted comprehensive ablation experiments examining joint versus separate training strategies and input feature configurations. Results in Table 16 demonstrate that joint optimization improves appliance state recognition F1-score by 6.44% over separate state-only training (88.27% 94.71%) and reduces injection estimation RMSE by 5.4% compared to separate injection-only training (0.9521 0.9008). The synergistic relationship between tasks enables the model to leverage accurate state predictions to refine injection estimates while using injection awareness to prevent false appliance detections during generation fluctuations.
Feature ablation studies (Figure 7) reveal that reactive power provides critical discriminative information. On laboratory data with random injection patterns, adding reactive power to active power improves F1-score by 12.23% (72.31% 84.54%), validating our insight that inverter near-unity power factor operation preserves appliance-specific reactive signatures. While external features such as weather data and time encoding provide marginal improvements on synthetic PV datasets, the core active and reactive power (P+Q) configuration achieves robust performance without communication overhead, enabling practical edge deployment. Complete ablation results and detailed analysis are provided in Appendix F.
6.4. Key Insights and Analysis
@@
Traditional methods, such as FHMM, fail due to violated mathematical assumptions, and their factorial structure assumes discrete states incompatible with continuous injection variations. Notably, XGBoost serves as a strong baseline that achieves competitive performance (74.41% average F1 on REDD from Table 12) through implicit feature engineering via recursive tree splits on the dual-channel input. However, it is fundamentally limited by its inability to provide explicit injection estimates and its reliance on pattern memorization, which makes it susceptible to distributional shifts.
@@ 7. Conclusion
Our findings underscore DualNILM’s potential for robust NILM in modern energy systems that integrate renewable energy sources, and our provided open research infrastructure [link] will help the NILM community continue to advance research in this critical direction.
Appendix
Ethical Considerations
This research adheres to established ethical principles for computational research and poses minimal ethical risks. Our study involves purely technical methodology development for energy disaggregation systems using laboratory-controlled measurements and synthetic datasets derived from publicly available benchmarks (REDD, UK-DALE).
Data Sources and Privacy. All experimental data consists of: (1) controlled laboratory measurements from our own micro-inverter setup, and (2) synthetic photovoltaic injection patterns generated using established solar modeling techniques without personal information. No human subjects, user behavioral data, or personally identifiable information were involved in this research.
Societal Impact. The proposed DualNILM framework aims to enhance energy monitoring capabilities in renewable-integrated smart grids, contributing to more efficient energy management and supporting the transition to sustainable energy systems. The method addresses a purely technical challenge without political, religious, or culturally sensitive implications.
Open Science. To promote reproducibility and advance community research, we provide open access to our synthetic PV-augmented datasets and benchmark implementations, following best practices for responsible data sharing in the energy informatics community.
@@ Appendix A Theoretical Analysis of the Challenge: When Negative Flows Disrupt Classical NILM
A.1. Matrix Formulation of NILM
@@
A.2. Why BTM Injection Complicates NILM
In the language of signal processing and blind source separation (BSS), classical NILM can be viewed as a single-channel mixture of purely nonnegative sources. The geometry of such mixtures is conic: if each device’s load lies in the nonnegative orthant, standard factorization approaches (e.g., nonnegative matrix factorization) can leverage convexity to help isolate distinct appliance patterns (Lee and Seung, 1999).
@@ A.2. Why BTM Injection Complicates NILM
A.3. Relaxation to Negative Flows and Identifiability
@@
Degraded Identifiability and Solution Stability. Under classical NILM assumptions (fully nonnegative), more stable decompositions can sometimes be achieved when the underlying sources have distinct support or exhibit certain ”sparse plus low-rank” structures. These identifiability arguments are often reminiscent of theorems in blind source separation (BSS) and NMF, where the nonnegative cone enforces a geometric structure that ”pins down” solutions. Once negative values are allowed (even for a single source), the problem becomes closer to an unconstrained BSS, for which solution identifiability is well known to deteriorate significantly. In more formal terms, NMF-based identifiability relies on the fact that data points lie in a convex cone generated by the basis columns (Lee and Seung, 1999); introducing negative components breaks those convex cone arguments.
A.4. Signal Space Interpretation and Rank Constraints